OpenStudy (andijo76):

use a normal approximation to find the probability of the indicated number of voters. in this case assume that 145 eligible voters aged 18-24 are randomly selected. Suppose a previous study showed that among eligible voters aged 18-24, 22% of them voted The probability that fewer than 35 of 145 eligible voters voted is what?

OpenStudy (andijo76):

the answer i have is this

OpenStudy (andijo76):

n=145 p=0.22 mean, mu=np=31.9 SD=145*.22*(1-.22)=4.9881 z=(35-31.9)/4.9881=0.6214 P(less than 37 voted)=P(z is less than .6215 now i am stuck